OpenStudy (anonymous):

Check my answer? I have to rewrite f(x)= 2x^2 + 5x + 3 in general form. This is what I got, am i right? 2. f(x)= 2x^2 + 5x + 3 fx)= 2x^2+5x+3 f(x)= 2x2 +2x +3x +3 f(x) = 2x(x + 1) + 3(x + 1) f(x) = (2x + 3)(x+1)

jimthompson5910 (jim_thompson5910):

I think they want it in 4p(y - k) = (x - h)^2 form

jimthompson5910 (jim_thompson5910):

where (h,k) is the vertex and p is the focal distance

OpenStudy (anonymous):

explain?

jimthompson5910 (jim_thompson5910):

have you seen vertex form before?

OpenStudy (anonymous):

yes, but i dont remember what itlooks like

jimthompson5910 (jim_thompson5910):

vertex form is y = a(x-h)^2 + k where (h,k) is the vertex and 'a' determines the shape and direction the parabola opens up

jimthompson5910 (jim_thompson5910):

so what do you get when you convert the given equation into vertex form?

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

2x^2 + 5x + 3 is in the form ax^2 + bx + c in this case a = 2 b = 5 c = 3

jimthompson5910 (jim_thompson5910):

the axis of symmetry is found by this formula x = -b/(2a)

jimthompson5910 (jim_thompson5910):

so what is the axis of symmetry in this case?

OpenStudy (anonymous):

-5/(2)(2)

OpenStudy (anonymous):

-5/4

jimthompson5910 (jim_thompson5910):

that's the x coordinate of the vertex, so h = -5/4 plug x = -5/4 into the original equation, and evaluate to get the y coordinate of the vertex

OpenStudy (anonymous):

2(-5/4)^2 + 5(-5/4) + 3

jimthompson5910 (jim_thompson5910):

good, keep going

OpenStudy (anonymous):

idk how to multiply fractions and whole numbers honestly

jimthompson5910 (jim_thompson5910):

think of whole numbers as that number over 1 example: 2 = 2/1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

-5/2x^2 - 25/4 +3

jimthompson5910 (jim_thompson5910):

remember you replaced x with -5/4

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

-5/2^2 - 25/4 + 3

jimthompson5910 (jim_thompson5910):

2(-5/4)^2 + 5(-5/4) + 3 2(25/16) + 5(-5/4) + 3 50/16 -25/4 + 3 25/8 - 25/4 + 3 I'll let you finish this piece up

OpenStudy (anonymous):

woah you lost me

jimthompson5910 (jim_thompson5910):

sorry, basically I'm following PEMDAS 2(-5/4)^2 + 5(-5/4) + 3 .. plug in x = -5/4 2(25/16) + 5(-5/4) + 3 .. exponents (square-5/4 to get 25/4) (2/1)*(25/16) + 5(-5/4) + 3 .. change 2 to 2/1 50/16 -25/4 + 3 ... multiplication 25/8 - 25/4 + 3 ... reducing 50/15 to 25/8

jimthompson5910 (jim_thompson5910):

hopefully that's a bit clearer